How fast must an 816-kg Volkswagen travel to have the same momentum as (a) a 2650-kg Cadillac going 16.0 km/h? (b) a 9080-kg truck also going 16.0 km/hr?

Grade:11

2 Answers

erra akhil
8 years ago
a)52 kmph approx. b)178kmph
M1V1=M2V2
Aditi Chauhan
8 years ago
(a) Mass of Volkswagen, m = 816-kg
Mass of Cadillac, m1 = 2650-kg
Velocity of Cadillac, v1 = 16.0 km/h
Let us consider velocity of the Volkswagen is v.
Momentum of a particle (p) is equal to the mass of the particle (m) times velocity of the particle (v).
So the momentum of Volkswagen (p) will be,
p = mv
= (816-kg) v …… (1)
and
the momentum of Cadillac (p1) will be,
p1 = m1 v1
= (2650-kg) (16.0 km/h)
= 42400 kg. km/h …… (2)
Since the momentum of the Volkswagen and Cadillac is same, therefore,
p = p1
(816-kg) v = 42400 kg. km/h
So, v = (42400 kg. km/h)/ 816-kg
= 51.96 km/h …… (3)
Rounding off to three significant figures the velocity of the Cadillac will be 52.0 km/h.
(b) Mass of Volkswagen, m = 816-kg
Mass of truck, m2 = 9080-kg
Velocity of truck, v2 = 16.0 km/h
Let us consider velocity of the Volkswagen is v.
Momentum of a particle (p) is equal to the mass of the particle (m) times velocity of the particle (v).
So the momentum of Volkswagen (p) will be,
p = mv
= (816-kg) v …… (4)
and
the momentum of truck (p2) will be,
p2 = m2 v2
= (9080-kg) (16.0 km/h)
= 145280 kg. km/h …… (5)
Since the momentum of the Volkswagen and truck is same, therefore,
p = p2
(816-kg) v = 145280 kg. km/h
So, v = (145280 kg. km/h)/ 816-kg
= 178.04 km/h …… (6)
Rounding off to three significant figures the velocity of the truck will be 178.04 km/h.